OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 8) and (12, 1)? A.-9/7 B.-7/9 C.7/9 D.9/7 @myininaya @beccaboo333 @destinymasha @amistre64 @hartnn @sam.

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

now im vconfused

hartnn (hartnn):

x1,y1 = 3,8 x2,y2 =12,1 lets first take the difference in y co-ordinates (the numerator) \(y_1 - y_2 = 8 -1 = ...\)

OpenStudy (anonymous):

7

OpenStudy (anonymous):

then what

hartnn (hartnn):

yes, now denominator \(x_1 -x_2 = 3-12 = ...?\)

OpenStudy (anonymous):

-9?

hartnn (hartnn):

correct! so your slope m = 7/(-9) = -7/9 :)

OpenStudy (anonymous):

so it would be b?

hartnn (hartnn):

yes!